OpenStudy (anonymous):

Which of the ordered pairs is a solution of the system of equations? y = -3/5x x - 3y = -42

OpenStudy (anonymous):

(-0.0429,13.9857) AND (-41.957, 0.0143)

OpenStudy (anonymous):

i dont get this answer

OpenStudy (anonymous):

??

OpenStudy (anonymous):

y = -3/5x --------------------------eq(1) x - 3y = -42-------------------------eq(2) (1)->(2): x-3(-3/5x)=-42 5x^2 -210x+9=0 x=-0.042901 , x=-41.9571 Substitute those x into eq(1) to find y.